can u help me with this question?
the length of a rectangle is five more than four times the width. the perimeter of the rectangle is 140 inches. find the dimensions of the rectangle

Answer by stanbon(75887)

the length of a rectangle is five more than four times the width. the perimeter of the rectangle is 140 inches. find the dimensions of the rectangle
------
width = x
length = 4x+5
----
Perimeter = 2(length + width)
140 = 2(4x+5 + x)
70 = 5x + 5
5x = 65
x = 13 (width)
4x+5 = 4*13+5 = 57 (length)
==============================
Cheers,
Stan H.
------------------------------